Two angles are supplements. The larger angle is 30 more than three times the smaller. Find the measures of these two angles.
Found 2 solutions by oberobic, stanbon:
Answer by oberobic(2304)

x + y = 180
x = 3y + 30
substitute
3y+30 +y = 180
4y = 150
y = 37.5 degrees
x = 3*37.5+30
x = 142.5 degrees
.
Done.

Answer by stanbon(75887)

Two angles are supplements. The larger angle is 30 more than three times the smaller. Find the measures of these two angles.
---
Which one you designate as larger or smaller is arbitrary.
1st: x (smaller)
2nd: 180-x (larger)
------------------------
Equation:
180-x = 3x+30
150 = 4x
---
x = 37.5 degrees
180-x = 142.5 degrees
=======================
Cheers,
Stan H.
==============